OpenStudy (anonymous):

How do you solve exponential equations that have different bases, such as 6^x-1 = 4.5^ 3x-1

OpenStudy (anonymous):

Hint :take log to both sides to get the power down .

OpenStudy (anonymous):

Assuming \[6^{x-1}=4.5^{3x-1}\] You have to use logs. log both sides: \[\log(6^{x-1})=\log(4.5^{3x-1})\] Use the power rule to move the exponents as coefficients. \[(x-1)\log(6)=(3x-1)\log(4.5)\] Now distribute the coefficients to both sides, combine like terms, and move x's to one side. Then you can factor out the x and divide for the answer.

OpenStudy (anonymous):

Thank you for your insight! I just completely drew a blank on this one.